# PARCO: Learning Parallel Autoregressive Policies for Efficient Multi-Agent Combinatorial Optimization

### v1

@@ 4.2 Parallel Autoregressive Model
PARCO additionally defines an agent projection layer that maps agents into the same embedding space as the problem nodes, whose exact definition heavily depends on the problem. In routing problems, for example, agent embeddings can be represented by the depot location they are starting from and their capacities, while an agent in scheduling problems can be represented by the machine whose schedule it is constructing. Lastly, encoded actions and agents can further be enhanced by performing cross-attention between them, depending on the structure of the problem (Kwon et al., 2021). We describe the specific encoder details for the different problems covered in this work in the Appendix.
Having a separate agent encoder contrasts other multi-agent (NCO) approaches like Zhang et al. (2020b), which use separate policies $\pi^{k}_{\theta_{m}}$ for each agent. This has a substantial drawback, as a model trained for a given number of agents cannot be used to solve problems with a different number of agents. The PARCO encoder, on the other hand, can generate representations for arbitrary agents, making it agnostic to the problem size.
Decoder
The decoder is the core of the PARCO architecture. We adopt the pointer mechanism from Kool et al. (2018), which yields unnormalized log-probabilities $\bm{z}\in\mathbb{R}^{N}$ for each action/node:
@@
$\bm{z}=C\cdot tanh\left(\frac{\bm{u}^{\top}L}{\sqrt{d_{h}}}\right)$ (3)
where $C$ is a scale parameter, $L$ is a projection of the node embeddings $\bm{h}$ and $\bm{u}$ is the output of an MHA layer, in which the query is defined by a single context embedding $\bm{q}^{c}_{t}\in\mathbb{R}^{d_{h}}$, specifying the current state of the problem at decoding step $t$ and keys and values are the node embeddings from the encoder. Each attention head of the MHA layer can further fuse the projection of the encoder embeddings with dynamic embeddings representing current information about the particular node, i.e.
$\text{Attn}(\bm{q}^{c}_{t}W^{Q}_{i},\,\bm{h}W^{K}_{i}+W^{K}_{\delta}\delta_{t},\,\bm{h}W^{V}_{i}+W^{V}_{\delta}\delta_{t})$ (4)
with $\delta_{t}$ encapsulating the dynamic elements of the problem nodes and $W^{K}_{\delta},W^{V}_{\delta}\in\mathbb{R}^{|\delta|\times d_{m}}$ projecting them into embedding space.
For PARCO we introduce a novel Multiple Pointer Mechanism based on Eq. 3 of the following form:
$Z=C\cdot tanh\left(\frac{UL^{\top}}{\sqrt{d_{h}}}\right)$ (5)
Here, $U\in\mathbb{R}^{m\times d_{h}}$ is the output of an equivalent MHA layer than the one used in Kool et al. (2018). However, instead of a single context embedding, PARCO uses the $m$ agent embeddings of the encoder, enriched by the current state’s context and individual agent’s state embeddings, as queries $Q^{c}_{t}\in\mathbb{R}^{m\times d_{h}}$. The resulting $Z\in\mathbb{R}^{m\times N}$ log-probabilities are used to sample $m$ actions - one for each agent - in parallel.
Communication Layers
To enable effective agent coordination, we introduce a communication layer before the Multiple Pointer Mechanism, which updates the agent embeddings through messages exchanged with other agents. This component is vital for parallel decoding of multiple actions, as the quality of an action chosen by one agent is significantly influenced by how it affects the actions of other agents.
Given the $m$ agent queries at decoding step $t$, the communication layer applies a transformer block (Vaswani et al., 2017) to capture intra-agent relationships and spot potential conflicts between agents:
$H^{\prime}$ $=\text{Norm}(\text{MHA}(Q^{c}_{t},Q^{c}_{t},Q^{c}_{t})+Q^{c}_{t})$ (6)
$Q^{\prime}_{t}$ $=\text{Norm}(\text{MLP}(H^{\prime})+H^{\prime})$ (7)
where Norm denotes a normalization layer (Ioffe and Szegedy, 2015; Zhang and Sennrich, 2019) and MLP represents the multi-layer perceptron. Self-attention in the MHA layer enables message passing between agents based on their embeddings. By applying the communication layer after fusing the agent embeddings with the dynamic context embeddings, it can capture dynamic relationships that could not be resolved during encoding, thereby enabling the decoder to make more informed decisions. The updated query $Q^{\prime}_{s}$ is finally passed into the Multiple Pointer Mechanism.
4.2.1 Conflicts Handlers
When sampling from the probability distribution generated by the Multiple Pointer Mechanism, it is possible for multiple agents to select the same node simultaneously, resulting in a conflict. These conflicts must be resolved to ensure the generation of a feasible solution. To address this, we introduce a Priority-based Conflict Handler $\mathcal{H}$, which resolves such conflicts by leveraging a predefined priority scheme. The conflict handler can be defined as a function $\mathcal{H}:\mathbb{N}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{N}^{m}$ with number of agents $m$. The priorities are determined based on the current actions and states. Given an input vector of actions $\bm{a}=(a_{1},a_{2},\dots,a_{m})$ where $a_{i}\in\mathbb{N}$, the corresponding priority vector $\bm{p}=(p_{1},p_{2},\dots,p_{m})$ where $p_{i}\in\mathbb{R}$ and the fallback actions (i.e., the previous node) $\bm{r}=(r_{1},r_{2},\dots,r_{m})$ where $r_{i}\in\mathbb{N}$, the output is another vector of actions $\bm{a}^{\prime}=(a^{\prime}_{1},a^{\prime}_{2},\dots,a^{\prime}_{m})$ such that any conflicts are resolved based on the priority order as illustrated in Algorithm 1.
@@ 4.2.1 Conflicts Handlers
0:  Actions $\bm{a}\in\mathbb{N}^{m}$, Priorities $\bm{p}\in\mathbb{R}^{m}$, Fallback actions $\bm{r}\in\mathbb{N}^{m}$
0:  Resolved Actions $\bm{a^{\prime}}\in\mathbb{N}^{m}$
1:  $\sigma\leftarrow\text{argsort}(\bm{p},\text{descending}=\text{True})$Sort indices by priority}
2:  $\hat{\bm{a}}\leftarrow\bm{a}[\sigma]$Reorder actions according to priority}
3:  Initialize conflict mask $M\leftarrow\bm{0}^{m}$
@@
8:  end for
9:  $\hat{\bm{a}}\leftarrow(1-M)\odot\hat{\bm{a}}+M\cdot r$Resolve conflicts by assigning fallback action $r$}
10:  $\bm{a}^{\prime}\leftarrow\hat{\bm{a}}[\sigma^{-1}]$Reorder actions back to original sequence}
11:  return  $\bm{a^{\prime}}$
4.2.2 Step Definition
We define a single parallel step of the model similarly to Eq. 1 as follows:
$p_{\theta}(\bm{a}|\bm{x})$ $=\prod_{i=1}^{m}g_{\theta}(a_{t,i}^{\prime}|a_{t-1,i}^{\prime},\dots,a_{0,i}^{\prime},\bm{h})$ (8)
Here, $a_{t,i}^{\prime}$ denotes the action executed by agent $i$ at time $t$, after passing through the conflict handler $\mathcal{H}$:
$a_{t,i}^{\prime}=\mathcal{H}(a_{t,1},a_{t,2},\dots,a_{t,m};\bm{p}_{t},\bm{r}_{t}),$
where $\bm{p}_{t}$ and $\bm{r}_{t}$ are priorities and fallback actions, respectively, at the current decoding step $t$. The conflict handler $\mathcal{H}$ ensures that multiple agents do not select the same node, thus maintaining the integrity and feasibility of the solution.
4.3 Training
PARCO is a centralized multi-agent sequential decision-making model, with a shared policy $\pi_{\theta}$ for all agents and a global reward $R$. Thus, PARCO can be trained using any of the training algorithms adopted in the single-agent NCO literature. We train PARCO via the REINFORCE gradient estimator (Williams, 1992) with a shared baseline as outlined by Kwon et al. (2020) and Kim et al. (2022):
$\nabla_{\theta}\mathcal{L}\approx\frac{1}{B\cdot L}\sum_{i=1}^{B}\sum_{j=1}^{L}G_{ij}\nabla_{\theta}\,\text{log}\,p_{\theta}(A_{ij}|\bm{x}_{i})$ (9)
where $B$ is the size of the mini-batch and $G_{ij}$ is the advantage $R(A_{ij},\bm{x}_{i})-b^{\text{shared}}(\bm{x}_{i})$ of a solution $A_{ij}$ compared to the shared baseline $b^{\text{shared}}_{i}$ of problem instance $\bm{x}_{i}$.
@@ 5 Experiments
In this section, we present the experimental results of PARCO in two routing problems, the min-max heterogenous capacitated vehicle routing problem (HCVRP) and the open multi-depot capacitated pickup and delivery problem (OMDCPDP), and a scheduling problem, namely the flexible flow shop problem (FFSP). We provide more details about the problem and experimental setups in Appendix A and Appendix B respectively.
5.1 Problem Descriptions
@@
The min-max HCVRP consists of $m$ agents sequentially visiting customers to satisfy their demands, with constraints including each customer can be visited exactly once and the amount of demand satisfied by a single vehicle in a trip cannot exceed its capacity, which can be reloaded by going back to the depot. The goal is to minimize the makespan, i.e., the worst route. Baselines include SISR (Christiaens and Vanden Berghe, 2020), Genetic Algorithm (GA) (Karakatič and Podgorelec, 2015), Simulated Annealing (SA) (İlhan, 2021), the Attention Model (AM) (Kool et al., 2018), Equity Transformer (ET) (Son et al., 2024), the model from Li et al. (2022) (DRLLi), and the state-of-the-art neural baseline 2D-Ptr (Liu et al., 2024c).
@@
The OMDCPDP problem is a practical variant of the pickup and delivery problem in which agents have a stacking limit of orders that can be carried at any given time. Pickup and delivery locations are paired, and pickups must be visited before deliveries. Multiple agents start from different depots with no need to go back (open). The goal is to minimize the sum of arrival times to delivery locations, i.e. minimizing the cumulative lateness. We include ORTools (Furnon and Perron, 2024) as a classical baseline, the Heterogeneous Attention Model (HAM) (Li et al., 2021a) for sequential decision-making and MAPDP (Zong et al., 2022) for parallel decision-making.
@@
In the flexible flow shop problem (FFSP), $N$ jobs must be processed across $S$ stages, each with multiple machines ($m>1$). Jobs follow a specified sequence through these stages, but within each stage, any available machine can process the job, with the key constraint that no machine can handle more than one job simultaneously. The goal is to schedule the jobs so that all jobs are finished in the shortest time possible. Notable benchmarks include the MatNet model (Kwon et al., 2021), the Random and Shortest Job First (SJF) dispatching rules, as well as the evolutionary algorithms Particle Swarm Optimization (PSO), and Genetic Algorithm (GA) (Hejazi and Saghafian, 2005).
Table 1: Benchmarks and results of our model for min-max HCVRP problems of varying sizes and agent numbers. Highlighting cost ($\downarrow$) and average gaps ($\downarrow$) from the best-known solutions of classical heuristic solvers. Inference times are shown in seconds in parentheses ()($\cdot$)( ⋅ ). (g.) refers to greedy performance while (s.) refers to sampling 1280 solutions.
[…]
$m$ 3 5 7 3 5 7 3 5 7 avg.
SISR 6.57 (271) 4.00 (274) 2.91 (276) 8.52 (425) 5.10(430) 3.69 (434) 10.29 (615) 6.17 (623) 4.45 (625) 0.00
GA 9.21 (233) 6.89 (320) 5.98 (405) 12.32 (347) 8.95 (465) 7.58 (578) 15.33 (479) 10.93 (623) 9.10 (772) 74.90
SA 7.04 (130) 4.39 (289) 3.30 (362) 9.17 (318) 5.61 (417) 4.17 (515) 11.13 (434) 6.80 (557) 5.01 (678) 10.21
AM (g.) 8.49 (0.08) 5.51 (0.08) 4.15 (0.09) 10.81 (0.10) 6.87 (0.11) 5.18 (0.10) 12.68 (0.14) 8.10 (0.13) 6.13 (0.13) 33.80
ET (g.) 7.58 (0.15) 4.76 (0.17) 3.58 (0.16) 9.76 (0.21) 6.01 (0.20) 4.43 (0.23) 11.74 (0.25) 7.25 (0.25) 5.23 (0.26) 17.67
DRLLi (g.) 7.43 (0.19) 4.71 (0.22) 3.60 (0.25) 9.64 (0.25) 5.97 (0.30) 4.52 (0.33) 11.44 (0.32) 7.06 (0.37) 5.38 (0.43) 17.08
2D-Ptr (g.) 7.20 (0.11) 4.48 (0.11) 3.31 (0.11) 9.24 (0.15) 5.65 (0.15) 4.14 (0.14) 11.12 (0.18) 6.75 (0.18) 4.92 (0.17) 10.54
PARCO (g.) 7.12 (0.12) 4.40 (0.11) 3.25 (0.11) 9.14 (0.15) 5.53 (0.15) 4.04 (0.15) 10.98 (0.19) 6.61 (0.18) 4.79 (0.18) 8.56
AM (s.) 7.62 (0.14) 4.82 (0.13) 3.63 (0.14) 9.92 (0.20) 6.19 (0.21) 4.64 (0.22) 11.82 (0.29) 7.45 (0.28) 5.58 (0.28) 20.69
ET (s.) 7.14 (0.21) 4.46 (0.22) 3.33 (0.22) 9.19 (0.30) 5.64 (0.30) 4.17 (0.31) 11.20 (0.41) 6.85 (0.38) 4.98 (0.40) 10.89
DRLLi (s.) 6.97 (0.30) 4.34 (0.36) 3.25 (0.43) 9.10 (0.45) 5.54 (0.55) 4.13 (0.65) 10.90 (0.60) 6.65 (0.76) 4.98 (0.92) 8.78
2D-Ptr (s.) 6.82 (0.13) 4.20 (0.13) 3.09 (0.14) 8.85 (0.17) 5.36 (0.18) 3.90 (0.19) 10.71 (0.22) 6.46 (0.23) 4.68 (0.24) 4.84
PARCO (s.) 6.82 (0.31) 4.17 (0.31) 3.06 (0.31) 8.79 (0.36) 5.28 (0.36) 3.83 (0.35) 10.61 (0.40) 6.36 (0.40) 4.58 (0.40) 3.65
5.2 Experimental Setup
We perform all experiments on a machine equipped with two Intel(R) Xeon(R) Gold 6338 CPU @ 2.00GHz CPUs with a total $128$ threads and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM. Training runs of PARCO take less than 24 hours each. During inference, we employ only one CPU and a single GPU. We report key metrics such as solution cost, inference times, and gaps in best-known solutions. We keep most settings of PARCO consistent across experiments, i.e., we use a Communication Layer and Priority-based Conflict Handling based on the highest probability action from the model output. We provide additional details regarding training and testing setups in the Appendix.
(a) Effect of Communication Layers.
(b) Effect of Conflict Handlers.
(c) PARCO vs AR decoding speed.
Figure 3: Ablation studies on PARCO components. PARCO scales better than AR models as ET.
Table 2: Benchmarks and results of our model for OMDCPDP problems of varying sizes and agent numbers.
[…]
$N$ 50 100 200 1000 avg.
$m$ 10 18 25 20 35 50 40 70 100 100 250 500
OR-Tools 23.73 18.64 16.92 43.34 34.80 31.89 78.35 64.93 60.16 562.11 333.01 310.55 0.00%
HAM (g.) 34.85 24.79 18.52 65.95 51.56 35.41 124.16 98.32 67.45 740.34 475.76 326.07 33.52%
MAPDP (g.) 27.95 20.13 17.54 52.44 38.23 $3$ 10.80%
PARCO (g.) 24.70 19.19 17.44 44.88 35.51 32.79 81.14 66.95 61.95 531.95 344.67 302.09 1.96%
HAM (s.) 32.84 25.45 18.12 63.59 49.12 34.90 122.77 96.32 67.01 739.77 471.09 321.50 31.02%
MAPDP (s.) 25.49 20.02 17.17 49.55 37.37 33.25 - - - - - - 7.04%
PARCO (s.) 24.27 18.71 17.00 43.36 34.82 32.19 79.23 65.83 61.12 526.52 338.78 298.71 -0.01%
Table 3: FFSP results. PARCO improves the MatNet performance and is also more than 4×4$\times$4 × faster overall.
Method FFSP20 FFSP50 FFSP100
MS Gap Time (training) MS Gap Time (training) MS Gap Time (training)
CPLEX 46.4 21.0 60s - -
CPLEX 36.6 6.4 600s - -
Random 47.8 22.9 0.1s 93.2 44 0.2s 167.2 78.0 0.3s
Shortest Job First 31.3 6.4 0.1s 57.0 7.8 0.1s 99.3 10.1 0.2s
Genetic Algorithm 30.6 5.7 25.4s 56.4 7.2 57.8s 98.7 9.5 104.5s
Particle Swarm Opt. 29.1 4.2 46.8s 55.1 5.9 92.9s 97.3 8.1 173.4s
MatNet 27.26 3.0 0.9s (2h) 51.52 0.5 2.1s (5.25h) 91.58 0.0 4.9s (28h)
PARCO (w/o CL) 27.02 2.1 0.2s (0.25h) 52.29 2.0 0.4s (1.25h) 92.15 2.0 0.9s (5.25h)
PARCO 26.47 0.0 0.2s (0.5h) 51.25 0.0 0.5s (2h) 91.55 0.0 1.1s (6.25h)

### v2

@@ 4.5 Decoder with Multiple Pointer Mechanism
where $\xi_{t}\in\mathbb{R}^{N\times k_{\xi}}$ are dynamic node features which are projected via ${\mathbf{W}}^{K}_{\xi},{\mathbf{W}}^{V}_{\xi}\in\mathbb{R}^{k_{\xi}\times d}$ for keys and values of the MHA, respectively. Further, $\mathbf{M}_{t}\in\mathbb{R}^{M\times N}$ is the current action mask at step $t$, avoiding agent representations to attend to infeasible actions.
We then obtain a joint logit space $\bm{u}$ across all agents:
@@
$\bm{u}={\beta}\cdot\text{tanh}\left(\frac{\bm{q}(\bm{h}_{n}{\mathbf{W}}^{L}+\xi_{t}{\mathbf{W}}^{L}_{\xi})^{\top}}{\sqrt{d}}\right)$ (7)
with learnable parameters ${\mathbf{W}}^{L}\in\mathbb{R}^{d\times d}$, ${\mathbf{W}}^{L}_{\xi}\in\mathbb{R}^{k_{\xi}\times d}$ and ${\beta}$ is a scale parameter, set to $10$ following Bello et al. (2016) to enhance exploration. The output logits $\bm{u}\in\mathbb{R}^{M\times N}$ are masked by setting infeasible actions given mask $\textbf{M}_{t}$ to minus infinity.The joint probability distribution over all agent actions is then computed as:
$p(\bm{a}_{t}|\bm{a}_{<t},\bm{h})=\prod_{m=1}^{M}\frac{\exp(\bm{u}_{m,a_{t}^{m}})}{\sum_{j=1}^{N}\exp(\bm{u}_{m,j})}$ (8)
where $\bm{a}_{t}=(a_{t}^{1},\ldots,a_{t}^{M})$ represents the joint action across all agents at step $t$.
4.6 Conflict Handlers
When sampling from the probability distribution $\bm{p}$ generated by the Multiple Pointer Mechanism, multiple agents may select the same action simultaneously, which can result in an infeasible solution in several CO problems: for instance, in vehicle routing problems, usually only one agent is allowed to visit a customer node. Thus, it is essential to choose how to deal with such a situation effectively. Conflict handling (i.e., tie-breaking) can be achieved by allowing a single agent among a number of agents that are in conflict to continue with its action while others revert to fallback actions, e.g., staying in their current position. A naive approach consists of randomly selecting an agent to perform the new action (Zong et al., 2022). However, this can be suboptimal since it excludes inductive biases that can be leveraged, such as learned representations.
@@ 4.6 Conflict Handlers
0:  Actions $\bm{a}\in\mathbb{N}^{M}$, Priorities $\bm{p}\in\mathbb{R}^{M}$, Fallback actions $\bm{r}\in\mathbb{R}^{M}$
0:  Resolved Actions $\bm{a}^{\prime}\in\mathbb{N}^{M}$ // Sort indices based on priorities in descending order.
1:  $\sigma\leftarrow\text{argsort}(\bm{p},\text{descending}=\text{True})$ // Reorder actions according to priority.
2:  $\hat{\bm{a}}\leftarrow\bm{a}[\sigma]$ // Initialize conflict mask (all zeros).
3:  $C\leftarrow\bm{0}^{M}$ // Check for conflicts in reordered actions.
@@
8:  end for// Resolve conflicts by assigning fallback actions.
9:  $\hat{\bm{a}}\leftarrow(1-C)\odot\hat{\bm{a}}+C\odot\bm{r}$ // Reorder resolved actions back to original order.
10:  $\bm{a}^{\prime}\leftarrow\hat{\bm{a}}[\sigma^{-1}]$
In PARCO, we propose Priority-based Conflict Handlers that leverage priorities as a tie-breaking rule. Such priorities can be based on heuristics – such as giving priority to agents close to completion or whose action results in the smallest immediate cost – or on learned priorities. In the latter case, the probability values of the selected actions $p(\bm{a}_{t})$ serve as an indicator for prioritizing certain agents: the higher its value, the higher the priority learned by the model to have that agent win the tie-break. Algorithm 1 shows an efficient vectorized implementation of the Priority-based Conflict Handler algorithm. In practice, we augment the conflict handler $\psi$ from Section 4.2 with priorities $\bm{p}\coloneq p(\bm{a}_{t})$ and fallback actions $\bm{r}$, i.e. $\psi(\cdot)\coloneq\psi(\bm{a}_{t},\bm{p},\bm{r})$. Fallback actions in PARCO correspond to “do nothing” operations – maintaining an agent’s current position (routing) or keeping a machine idle (scheduling) – which may result in slightly more solution construction steps but do not affect the final solution $\bm{a}$. This approach effectively handles conflicts by allowing the affected agents to reconsider their choices given the actions of (preceding) agents in the next decoding step.
4.7 Training Scheme
PARCO is a centralized multi-agent decision-making framework and can thus be trained by RL algorithms proposed in the single-agent NCO literature. We train PARCO by employing the REINFORCE gradient estimator (Williams, 1992) with a shared baseline as outlined by Kwon et al. (2020) and Kim et al. (2022):
$\nabla_{\theta}\mathcal{L}\approx\frac{1}{B\cdot S}\sum_{i=1}^{B}\sum_{j=1}^{S}G_{ij}\nabla_{\theta}\,\text{log}\,p_{\theta}(\bm{a}_{ij}|\bm{x}_{i})$ (9)
where $B$ is the batch size, $S$ the number of shared baseline samples, and $G_{ij}=R(\bm{a}_{ij},\bm{x}_{i})-b^{\text{shared}}(\bm{x}_{i})$ is the advantage of a solution $\bm{a}_{ij}$ compared to the shared baseline $b^{\text{shared}}_{i}$ of problem instance $\bm{x}_{i}$.
@@ 5 Experiments
We assess the effectiveness of PARCO on representative multi-agent CO problems, spanning both routing and scheduling domains. Specifically, we evaluate its performance on two challenging routing problems – the min-max heterogeneous capacitated vehicle routing problem (HCVRP) and the open multi-depot capacitated pickup and delivery problem (OMDCPDP) – as well as a scheduling problem, the flexible flow shop problem (FFSP).
5.1 Experimental Settings
We experiment on a workstation equipped with 2 Intel(R) Xeon(R) Gold 6338 CPUs and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM each. Training runs of PARCO take less than 24 hours each. During inference, we employ only one CPU and a single GPU. Additional details are available throughout Appendix B.
@@ 5.1 Experimental Settings
Problem. The min-max HCVRP involves ${M}$ agents serving customer demands while adhering to heterogeneous vehicle capacity constraints. Each vehicle can replenish its load by returning to the depot. The objective is to minimize the longest route taken by any agent (min-max), ensuring balanced workload distribution. Traditional solvers. We include state-of-the-art SISRs (Christiaens & Vanden Berghe, 2020), Genetic Algorithm (GA) (Karakatič & Podgorelec, 2015) and Simulated Annealing (SA) (İlhan, 2021). Neural baselines. We evaluate the sequential planning baselines Attention Model (AM) (Kool et al., 2019a), Equity Transformer (ET) (Son et al., 2024) and Decoupling Partition and Navigation (DPN) (Zheng et al., 2024a), and autoregressive models with agent selection DRLLi (Li et al., 2022) and state-of-the-art learning method 2D-Ptr (Liu et al., 2024b). Additional problem and experimental details are available in Section A.1 and Section B.1, respectively.
@@
Problem. The OMDCPDP is a challenging problem arising in last-mile delivery settings where $M$ agents starting from different locations (i.e., multiple depots) must pick up and deliver parcels without returning to their starting point (i.e., open). Agents have a capacity constraint for orders that can be carried out as a stacking limit: a tour can include more pickups than the constraint, but the agent must deliver corresponding orders so that carrying capacity is freed. The goal is to minimize the lateness, i.e., the sum of delivery arrival times. Traditional solvers. We include the popular and efficient optimization suite Google OR-Tools (Furnon & Perron, 2024) as a classical baseline. Neural baselines. We evaluate models specializing in pickup and delivery problems, including the autoregressive Heterogeneous Attention Model (HAM) (Li et al., 2021a) for sequential planning and MAPDP (Zong et al., 2022) for parallel planning. Additional problem and experimental details are available in Section A.2 and Section B.2, respectively.
@@
Problem. In FFSP, $N$ jobs must be processed by $M$ machines divided equally in $S$ stages. Jobs follow a specified sequence through these stages. Within each, any available machine can process the job, with the key constraint that no machine can handle multiple jobs simultaneously. The goal is to schedule the jobs so that all jobs are finished in the shortest time possible. Traditional solvers. We incorporate the widely used and powerful Gurobi solver (Gurobi Optimization, LLC, 2024) as a baseline. Furthermore, we include dispatching rules Random and Shortest Job First (SJF), Particle Swarm Optimization (PSO) (Singh & Mahapatra, 2012) and Genetic Algorithm (GA) (Hejazi & Saghafian, 2005). Neural baselines. Notable benchmarks include the Matrix Encoding Network (MatNet) (Kwon et al., 2021) which demonstrates superior performance on FFSP. Additional problem and experimental details are available in Section A.3 and Section B.3, respectively.
Table 1: Main results with different configurations in terms of problem size $N$ and number of agents $M$. The lower the better ($\downarrow$).
HCVRP
$N$ 60 100
$M$ 3 5 7 3 5 7
Metric Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
SISRs 6.57 0.00% 271s 4.00 0.00% 274s 2.91 0.00% 276s 10.29 0.00% 615s 6.17 0.00% 623s 4.45 0.00% 625s
GA 9.21 40.18% 233s 6.89 72.25% 320s 5.98 105.50% 405s 15.33 48.98% 479s 10.93 77.15% 623s 9.10 104.49% 772s
SA 7.04 7.15% 130s 4.39 9.75% 289s 3.30 13.40% 362s 11.13 8.16% 434s 6.80 10.21% 557s 5.01 12.58% 678s
AM (g.) 8.49 29.22% 0.08s 5.51 37.75% 0.08s 4.15 42.61% 0.09s 12.68 23.23% 0.14s 8.10 31.28% 0.13s 6.13 37.75% 0.13s
ET (g.) 7.58 15.37% 0.15s 4.76 19.00% 0.17s 3.58 23.02% 0.16s 11.74 14.09% 0.25s 7.25 17.50% 0.25s 5.23 17.53% 0.26s
DPN (g.) 7.50 14.16% 0.18s 4.60 15.00% 0.19s 3.45 18.56% 0.26s 11.54 12.15% 0.30s 6.94 12.48% 0.40s 4.98 11.91% 0.43s
DRLLi (g.) 7.43 13.09% 0.19s 4.71 17.75% 0.22s 3.60 23.71% 0.25s 11.44 11.18% 0.32s 7.06 14.42% 0.37s 5.38 20.90% 0.43s
2D-Ptr (g.) 7.20 9.59% 0.11s 4.48 12.00% 0.11s 3.31 13.75% 0.11s 11.12 8.07% 0.18s 6.75 9.40% 0.18s 4.92 10.56% 0.17s
PARCO (g.) 7.12 8.37% 0.04s 4.40 10.00% 0.05s 3.25 11.68% 0.05s 10.98 6.71% 0.06s 6.61 7.13% 0.05s 4.79 7.64% 0.05s
AM (s.) 7.62 15.98% 0.14s 4.82 20.50% 0.13s 3.63 24.74% 0.14s 11.82 14.87% 0.29s 7.45 20.75% 0.28s 5.58 25.39% 0.28s
ET (s.) 7.14 8.68% 0.21s 4.46 11.50% 0.22s 3.33 14.43% 0.22s 11.20 8.84% 0.41s 6.85 11.02% 0.38s 4.98 11.91% 0.40s
DPN (s.) 7.08 7.76% 0.25s 4.35 8.75% 0.28s 3.20 9.97% 0.38s 11.04 7.29% 0.48s 6.66 7.94% 0.52s 4.79 7.64% 0.78s
DRLLi (s.) 6.97 6.09% 0.30s 4.34 8.50% 0.36s 3.25 11.68% 0.43s 10.90 5.93% 0.60s 6.65 7.78% 0.76s 4.98 11.91% 0.92s
2D-Ptr (s.) 6.82 3.81% 0.13s 4.20 5.00% 0.13s 3.09 6.19% 0.14s 10.71 4.08% 0.22s 6.46 4.70% 0.23s 4.68 5.17% 0.24s
PARCO (s.) 6.82 3.81% 0.05s 4.17 4.25% 0.05s 3.06 5.15% 0.07s 10.61 3.11% 0.08s 6.36 3.08% 0.08s 4.58 2.92% 0.09s
OMDCPDP
$N$ 50 100
$M$ 5 7 10 10 15 20
Metric Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
OR-Tools 37.61 4.61% 30s 30.18 3.38% 30s 24.48 1.79% 30s 66.78 4.99% 60s 51.49 3.15% 60s 43.90 1.53% 60s
HAM (g.) 39.67 10.30% 0.11s 31.49 7.85% 0.11s 29.24 21.23% 0.13s 71.12 11.60% 0.24s 54.31 8.69% 0.27s 53.73 23.93% 0.29s
MAPDP (g.) 37.36 4.09% 0.02s 30.36 4.08% 0.01s 24.88 3.46% 0.01s 66.54 4.71% 0.01s 52.08 4.34% 0.01s 44.71 3.40% 0.01s
PARCO (g.) 37.27 3.84% 0.02s 30.12 3.27% 0.01s 24.72 2.81% 0.01s 65.85 3.67% 0.02s 51.45 3.11% 0.01s 44.46 2.84% 0.01s
HAM (s.) 36.26 1.17% 1.18s 29.54 1.38% 1.31s 27.77 15.24% 1.37s 66.91 5.29% 2.33s 51.60 3.42% 2.53s 52.24 20.52% 2.72s
MAPDP (s.) 35.64 0.02% 0.02s 29.03 0.23% 0.02s 23.97 0.35% 0.01s 63.64 0.65% 0.03s 50.07 0.75% 0.03s 43.57 1.13% 0.02s
PARCO (s.) 35.64 0.00% 0.03s 28.96 0.00% 0.02s 23.89 0.00% 0.02s 63.20 0.00% 0.04s 49.69 0.00% 0.03s 43.08 0.00% 0.03s
FFSP
$N$ 20 50 100 50
$M$ 12 18 24 30
Metric Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
Gurobi (1m) 35.29 42.4% 60s - - 60s - - 60s - - 60s - - 60s - - 60s
Gurobi (10m) 31.61 27.6% 600s - - 600s - - 600s - - 600s - - 600s - - 600s
Random 47.89 93.3% 0.18s 93.34 89.4% 0.37s 167.22 86.9% 0.72s 67.03 112.1% 0.33s 54.48 130.9% 0.33s 46.84 137.0% 0.37s
SJF 31.27 26.2% 0.13s 56.94 15.6% 0.34s 99.27 11.0% 0.62s 38.01 20.3% 0.25s 29.39 24.6% 0.25s 24.62 24.6% 0.29s
GA 31.15 25.7% 21s 56.92 15.5% 44s 99.25 10.9% 89s 38.26 21.1% 47s 29.05 16.7% 50s 24.52 24.1% 55s
PSO 29.10 17.4% 46s 55.10 11.8% 82s 97.3 8.8% 154s 36.83 16.6% 85s 28.06 12.7% 89s 23.44 18.6% 95s
MatNet (g.) 27.26 10.0% 1.22s 51.52 4.6% 2.17s 91.58 2.4% 4.97s 34.82 10.2% 2.42s 27.52 16.7% 2.65s 23.65 19.7% 3.09s
PARCO (g.) 26.31 6.2% 0.26s 51.19 3.9% 0.52s 91.29 2.0% 0.89s 32.88 4.1% 0.50s 24.89 5.5% 0.44s 20.29 2.7% 0.41s
MatNet (s.) 25.44 2.7% 3.88s 49.68 0.8% 8.91s 89.72 0.3% 18s 33.45 5.9% 9.23s 26.00 10.2% 9.81s 22.51 13.9% 11s
PARCO (s.) 24.78 0.0% 0.99s 49.27 0.0% 1.97s 89.46 0.0% 4.04s 31.60 0.0% 1.89s 23.59 0.0% 1.68s 19.76 0.0% 1.54s
Figure 3: Analysis on PARCO components. (a) Full Communication Layers provide the best performance. (b) Average number of decoding steps required to construct solutions (bar heights) and training times (lines) for different models. (c) Random and Priority-based Conflict Handlers variants with heuristics and learned priorities with learned providing the best performance.